# ChartCoder: Advancing Multimodal Large Language Model for Chart-to-Code Generation

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in chart understanding tasks. However, interpreting charts with textual descriptions often leads to information loss, as it fails to fully capture the dense information embedded in charts. In contrast, parsing charts into code provides lossless representations that can effectively contain all critical details. Although existing open-source MLLMs have achieved success in chart understanding tasks, they still face two major challenges when applied to chart-to-code tasks.: (1) Low executability and poor restoration of chart details in the generated code and (2) Lack of large-scale and diverse training data. To address these challenges, we propose ChartCoder, the first dedicated chart-to-code MLLM, which leverages Code LLMs as the language backbone to enhance the executability of the generated code. Furthermore, we introduce Chart2Code-160k, the first large-scale and diverse dataset for chart-to-code generation, and propose the Snippet-of-Thought (SoT) method, which transforms direct chart-to-code generation data into step-by-step generation. Experiments demonstrate that ChartCoder, with only 7B parameters, surpasses existing open-source MLLMs on chart-to-code benchmarks, achieving superior chart restoration and code excitability. Our code will be available at https://github.com/thunlp/ChartCoder.
@@
Xuanle Zhao1, [^1] Equal contribution., Xianzhen Luo2, [^1] Equal contribution., Qi Shi1, [^2] Corresponding author., Chi Chen1, [^2] Corresponding author., Shuo Wang1, Wanxiang Che2, Zhiyuan Liu1, Maosong Sun1 1 Tsinghua University, Beijing, China 2 Harbin Institute of Technology, Harbin, China
@@ 1 Introduction
Recently, Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in addressing a wide range of visual tasks, such as captioning and question answering Zhang et al. (2024c); Wang et al. (2024b); Chen et al. (2023); Zeng et al. (2024). However, current models still face significant challenges in understanding and analyzing the dense visual information present in complex and informative images. As a significant form of information-intensive images, charts contain complex information such as data and structures, playing a pivotal role in effectively presenting details. The automation of chart comprehension and summarization has garnered significant attention from the research community.
@@
To advance chart understanding tasks, current studies leverage existing MLLMs and perform supervised fine-tuning (SFT) on various large-scale chart understanding datasets, such as chart question answering Methani et al. (2020) and chart-to-text generation Kantharaj et al. (2022), achieving state-of-the-art performance on existing chart understanding benchmarks.
@@
However, existing works generally treat charts as natural images and fine-tune models by generating natural language descriptions Zhang et al. (2024b); Han et al. (2023); Meng et al. (2024). This inevitably overlooks the dense information embedded within the charts, resulting in inefficiencies in analysis and comprehension. On the other hand, parsing a chart into code offers a lossless representation, providing a more efficient and comprehensive approach to understanding the chart by accurately capturing and summarizing all its information. Recent works Shi et al. (2024); Wu et al. (2024); Xia et al. (2024) have proposed various chart-to-code benchmarks, aiming to evaluate the chart reasoning abilities through code. However, current open-source MLLMs are not well-aligned with code generation tasks Zhang et al. (2024a), resulting in poor performance in parsing charts into corresponding code and limited execution rate of the generated code. As shown in Figure 1, the InternVL2-8B suffers from chart type errors and coordinate size mismatches when converting boxplots to code.
@@
To overcome the above challenges in chart-to-code generation, we first conduct an exploratory attempt by leveraging Code LLMs as the language backbone of the MLLMs and propose ChartCoder, the first dedicated chart-to-code MLLM, which incorporates a two-stage training paradigm that contains chart-to-text alignment and chart-to-code instruction tuning. However, compared to chart-to-text, the available chart-to-code dataset is significantly smaller in scale, making it insufficient to support effectively supervised fine-tuning. Therefore, to address the scarcity of data for the chart-to-code domain and train our proposed ChartCoder, we propose the first large-scale diverse and high-quality chart-to-code dataset named Chart2Code-160k along with the model, which contains 160k diverse chart-code pairs with 27 chart types. To enhance the model’s capacity to capture critical information, such as chart types and data values, and strengthen its reasoning ability, we propose the Snippet-of-Thought (SoT) method, which emphasizes critical information and optimizes the chart-to-code reasoning process. Specifically, we sample 50k chart-code pairs from the Chart2Code-160k, then utilize Chain-of-Thought (CoT) Wei et al. (2022) and Program-of-Thought (PoT) Chen et al. (2022) methods to extend direct generation to step-by-step generation, which aims to emphasize critical information in each step. Experimental results show that by utilizing our proposed Chart2Code-160k with the SoT method, ChartCoder, which, with only 7B parameters, outperforms all open-source MLLMs across various chart-to-code benchmarks. As shown in Figure 1, ChartCoder demonstrates a significantly higher ability to generate correct and executable code.
@@ 2.1 Chart Understanding
Chart understanding is a crucial area of research that encompasses both low-level tasks, such as extracting data from visual charts, and high-level tasks, including QA Masry et al. (2022), summarization Kantharaj et al. (2022), and re-rendering chart content Han et al. (2023). Previous approaches Singh et al. (2019); Methani et al. (2020) have typically relied on pipeline-based methods. However, these pipeline approaches often struggle with error accumulation across different stages, which limits their overall effectiveness and flexibility. Recent works have led to the development of end-to-end MLLMs Liu et al. (2023b, c); Han et al. (2023); Ye et al. (2023) specifically designed for chart-related tasks. Trained on extensive chart-specific datasets, these chart-domain MLLMs Xia et al. (2024); Zhang et al. (2024b); Masry et al. (2024); Yang et al. (2024b) have achieved superior performance across various chart-related tasks. However, existing studies typically describe charts in natural language, which inevitably overlooks the dense information embedded within them, leading to inefficiencies in analysis and understanding. In contrast, code serves as a lossless representation of charts, offering a more effective and expressive approach to capturing chart information, thereby facilitating the solution of various chart-related tasks.
@@ 2.2 MLLMs For Code
Multimodal code generation has recently garnered much more attention. Several works, such as MMCode Li et al. (2024b) and HumanEval-V Zhang et al. (2024a), have been developed to evaluate the capability of MLLMs in solving code problems that incorporate visual elements. Design2Code Si et al. (2024) and Web2Code Yun et al. (2024) evaluate the performance of MLLMs by focusing on code generation for HTML web page creation. Among the emerging tasks in this domain, chart-to-code generation has attracted significant interest as the visual elements of charts are more complex. This task challenges MLLMs to generate code that accurately reproduces a given chart or visual representation. Recent works like ChartMimic Shi et al. (2024) evaluate the reasoning ability of MLLMs in this context. Similarly, Plot2Code Wu et al. (2024) and ChartX Xia et al. (2024) also evaluate MLLMs code generation ability, especially for text and data reproducibility. To the best of our knowledge, no dedicated research has focused on solving the chart-to-code generation problem. Our work is the first to attempt to address this challenge.
@@ 3.1 Direct Chart-to-code Generation
Despite the availability of many chart reasoning instruction-tuning datasets, there is a notable lack of datasets specifically for chart-to-code tasks. Compared to chart reasoning data, chart-to-code data have the following distinct characteristics: (1) One-to-One Mapping: Unlike chart reasoning datasets, which could derive multiple question-answer pairs from a single chart, chart-to-code datasets require a one-to-one correspondence, demanding a large number of chart images for training. (2) Diversity Reflect on Charts: Unlike the diversity of chart reasoning data, which can be reflected in instructions, the diversity of chart-to-code data primarily lies in the variety of chart types and structures. (3) Syntax Constraints: Unlike the flexible natural language answers in chart reasoning tasks, the output code must strictly adhere to programming syntax to ensure executability.
@@
Specifically, we generate chart-to-code data through the following steps: First, we prompt the LLM to generate keywords within a specific domain and guide it to generate simulated data related to these keywords. Then, to ensure the diversity of chart types, we identify 27 commonly used chart types and manually write seed codes for each as in-context demonstrations. We further provide available functions such as plt.text() and parameters such as hatch=’/’ to encourage the generation of more diverse functions and parameters Xu et al. (2023), resulting in the chart structures more diversely. To enhance the generality of generated code, LLMs are encouraged to use standard libraries such as Matplotlib and Seaborn. Additionally, we explicitly define all parameters within the code itself, eliminating the need for external files such as CSVs. This ensures that the code can be executed directly and accurately to represent the chart details. The final step involved executing the generated code to produce the corresponding chart. We utilize the above process to yield 200k code snippets for charts. After executing the code and filtering out problematic charts, such as those with excessive pixels or ticks, we construct a high-quality dataset of 160k diverse chart-to-code pairs. These pairs are formatted as multimodal instruction-tuning samples in the unified structure of <chart image, input instruction, output code>.
@@ 3.3 Dataset Analysis
Chart2Code-160k dataset provides three key advantages: (1) The First Large-Scale Dataset: It contains 160k data pairs for instruction tuning, significantly surpassing the size of previous datasets. (2) Diverse Chart Structures and Types: It includes 27 different chart types, with diverse structures enabled by a wide variety of functions and parameters in the code. (3) Syntactically Correct and Executable Code: All corresponding code is syntactically correct and executable, with explicitly defined parameters that ensure precise alignment between chart structures and code representations. The comparisons of Chart2Code-160k with relevant chart-to-code datasets are listed in Table 1. Although numerous chart-related datasets have been proposed Liu et al. (2023a); Han et al. (2023), none have sufficiently addressed the challenge of chart-to-code tasks. Our proposed Chart2Code-160k fills this gap, equipping the model with advanced capabilities for downstream chart tasks.
@@ 4 ChartCoder Model
After constructing the Chart2Code-160k, we aimed to leverage the data to enhance the capacities of MLLMs to generate code from charts. Unlike previous methods that rely on general LLMs with a low proportion of code in their training corpus, which limits their coding capabilities, we pioneer the use of Code LLMs to enhance the coding abilities of MLLMs from scratch.
@@ 4.2 Model Training
Chart-to-text Alignment. The alignment process aims to endow the model with chart structure perception capability. In this stage, we freeze the language and vision encoder models and pre-train the vision-language connector Liu et al. (2023c). We collect and filter public chart corpora for alignment, which contains multiple tasks like chart caption and chart-to-table. Specifically, we use the following corpora: (1) UniChart Masry et al. (2023), (2) Chart-to-Text Kantharaj et al. (2022), (3) SciCap Hsu et al. (2021), and (4) SciCap+ Yang et al. (2024a). Additionally, we incorporate the LLaVA pre-training dataset Liu et al. (2023c) and our proposed Chart2Code-160k to achieve a more balanced coverage of concepts.
@@ 5.2 Main Results
Table 4 shows the low-level results. ChartCoder even slightly outperforms GPT-4o in layout, type and color, highlighting the diversity of our proposed Chart2Code-160k dataset. However, the text score of the ChartCoder is lower than GPT-4o, which is similar to the results of high-level scores. We believe this is due to the lack of specialized chart OCR-oriented training for our model. Nevertheless, our text accuracy still surpasses that of open-source models, indicating the effectiveness of our proposed ChartCoder model and Chart2Code-160k dataset. We further present some case studies on ChartMimic and compare ChartCoder with GPT-4o, Qwen2-VL-72B, and InternVL-8B. The results are shown in Figure 4, the outputs of ChartCoder are much more similar to the ground truth chart than open-source MLLMs.
@@ 5.3 Ablation Study
We perform extensive ablation experiments to validate the effectiveness of our proposed model and dataset. We divide the ablation study into three parts, and the results are shown in Table 5 (1) Code or general LLMs. To investigate whether employing Code LLMs as language backbone provides specific advantages in chart-to-code tasks and identify the nature of these potential benefits, we replace the Code LLM, DeepSeek Coder 6.7B, with general LLM, DeepSeek LLM 7B Bi et al. (2024), maintaining the same two-stage training procedures. The result shows that compared with general LLM, utilizing code LLM as the language backbone could significantly improve the execution rate, as well as the low-level and high-level scores. We further analyze the types of errors in the code that failed to execute and find that utilizing code LLMs significantly reduces syntax errors like missing closing quotation marks and type errors like incorrect argument type. (2) Resolution of vision encoders. Previous studies have indicated that performance on chart understanding tasks is resolution-dependent, with lower resolutions negatively impacting model performance Liu et al. (2024). To verify whether resolution affects chart-to-code tasks, we replace SigLIP-384 with CLIP-336 Radford et al. (2021) and maintain the other setting. The result shows that the resolution of the vision encoder generally does not affect the output code execution rate but slightly influences the high-level chart similarity. Through our analysis, we find that, similar to the challenges in chart understanding, this issue is caused by the negative impact of low resolution on the recognition of text and special symbols. However, as we utilize the Any Resolution strategy, this impact has been reduced significantly. (3) Dataset effectiveness. We design two scenarios to illustrate our proposed Chart2Code-160k dataset. Firstly, to evaluate our proposed SoT method to emphasize the critical information in the chart, we remove the 50k step-by-step generation data and train the model using only the direct generation data. The result shows it influences the low-level and high-level scores notably, especially in text content and data, which shows the role of emphasising critical information. Secondly, we select Qwen2-VL-7B as the baseline of open-source MLLM and directly fine-tune it on our proposed Chart2Code-160k datasets. The result illustrates that after fine-tuning, the performance improves significantly on all the metrics, demonstrating the effectiveness of Chart2Code-160k.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in chart understanding tasks. However, interpreting charts with textual descriptions often leads to information loss, as it fails to fully capture the dense information embedded in charts. In contrast, parsing charts into code provides lossless representations that can effectively contain all critical details. Although existing open-source MLLMs have achieved success in chart understanding tasks, they still face two major challenges when applied to chart-to-code tasks: (1) Low executability and poor restoration of chart details in the generated code and (2) Lack of large-scale and diverse training data. To address these challenges, we propose ChartCoder, the first dedicated chart-to-code MLLM, which leverages Code LLMs as the language backbone to enhance the executability of the generated code. Furthermore, we introduce Chart2Code-160k, the first large-scale and diverse dataset for chart-to-code generation, and propose the Snippet-of-Thought (SoT) method, which transforms direct chart-to-code generation data into step-by-step generation. Experiments demonstrate that ChartCoder, with only 7B parameters, surpasses existing open-source MLLMs on chart-to-code benchmarks, achieving superior chart restoration and code excitability. Our code is available at https://github.com/thunlp/ChartCoder.
@@
Xuanle Zhao1, [^1] Equal contribution., Xianzhen Luo2, [^1] Equal contribution., Qi Shi1, [^2] Corresponding author., Chi Chen1, [^2] Corresponding author., Shuo Wang1, Zhiyuan Liu1, Maosong Sun1 1 Tsinghua University, Beijing, China 2 Harbin Institute of Technology, Harbin, China 2429527z@gmail.com, xzluo@ir.hit.edu.cn
@@ 1 Introduction
Recently, Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in addressing a wide range of visual tasks, such as captioning and question answering Zhang et al. (2024c); Wang et al. (2024b); Bi et al. (2024b); Zhang et al. (2025). However, current models still face significant challenges in understanding and analyzing the dense visual information present in complex and informative images. As a significant form of information-intensive images, charts contain complex information such as data and structures, playing a pivotal role in effectively presenting details. The automation of chart comprehension and summarization has garnered significant attention from the research community.
@@
To advance chart understanding tasks, current studies leverage existing MLLMs and perform supervised fine-tuning (SFT) on various large-scale datasets, such as chart question answering Methani et al. (2020) and chart-to-text generation Kantharaj et al. (2022), achieving state-of-the-art performance on existing chart understanding benchmarks.
@@
However, existing works generally treat charts as natural images and fine-tune models by generating natural language descriptions Zhang et al. (2024b); Han et al. (2023); Meng et al. (2024); Bi et al. (2025). This inevitably overlooks the dense information embedded within the charts, resulting in inefficient analysis and comprehension. On the other hand, parsing a chart into code offers a lossless representation, providing a more efficient and comprehensive approach to understanding the chart by accurately capturing and summarizing all its information. Recent works Shi et al. (2024); Wu et al. (2024); Xia et al. (2024) have proposed various chart-to-code benchmarks, aiming to evaluate the chart reasoning abilities through code. However, current open-source MLLMs are not well-aligned with code generation tasks Zhang et al. (2024a), resulting in poor performance in parsing charts into corresponding code and limited execution rate of the generated code. Figure 1 shows that the InternVL2-8B suffers from chart type errors and coordinate size mismatches when converting boxplots to code.
@@
To overcome the above challenges in chart-to-code generation, we first conduct an exploratory attempt by leveraging Code LLMs as the language backbone of the MLLMs and propose ChartCoder, the first dedicated chart-to-code MLLM, which incorporates a two-stage training paradigm that contains chart-to-text alignment and chart-to-code instruction tuning. However, compared to chart-to-text, the available chart-to-code dataset is significantly smaller in scale, making it insufficient to support effectively supervised fine-tuning. Therefore, to address the scarcity of data for the chart-to-code domain and train our proposed ChartCoder, we propose the first large-scale, diverse and high-quality chart-to-code dataset named Chart2Code-160k along with the model, which contains 160k diverse chart-code pairs with 27 chart types. To enhance the model’s capacity to capture critical information, such as chart types and data values, and strengthen its reasoning ability, we propose the Snippet-of-Thought (SoT) method, which emphasizes critical information and optimizes the chart-to-code reasoning process. Specifically, we sample 50k chart-code pairs from the Chart2Code-160k, then utilize Chain-of-Thought (CoT) Wei et al. (2022) method to extend direct generation to step-by-step generation, which aims to emphasize critical information in each step. Experimental results show that by utilizing our proposed Chart2Code-160k with the SoT method, ChartCoder, which, with only 7B parameters, outperforms all open-source MLLMs across various chart-to-code benchmarks. As shown in Figure 1, ChartCoder demonstrates a significantly higher ability to generate correct and executable code.
@@ 2.1 Chart Understanding
Chart understanding is a crucial area of research that encompasses both low-level and high-level tasks. Previous approaches Singh et al. (2019); Methani et al. (2020) have typically relied on pipeline-based methods. However, these pipeline approaches often struggle with error accumulation across different stages, which limits their overall effectiveness and flexibility. Recent works have led to the development of end-to-end MLLMs Liu et al. (2023b, c); Yu et al. (2025) specifically designed for chart-related tasks. Trained on extensive chart-specific datasets, these chart-domain MLLMs Xia et al. (2024); Zhang et al. (2024b) have achieved superior performance across various chart-related tasks. However, existing studies typically describe charts in natural language, which inevitably overlooks the dense information embedded within them, leading to inefficiencies in analysis and understanding. In contrast, code serves as a lossless representation of charts, offering a more effective and expressive approach to capturing chart information, thereby facilitating the solution of various chart-related tasks.
@@ 2.2 MLLMs For Code
Multimodal code generation has recently garnered much more attention. Several works, such as MMCode Li et al. (2024b) and HumanEval-V Zhang et al. (2024a), have been developed to evaluate the capability of MLLMs in solving code problems that incorporate visual elements. Design2Code Si et al. (2024) and Web2Code Yun et al. (2024) evaluate the performance of MLLMs by focusing on code generation for HTML web page creation. Among the emerging tasks in this domain, chart-to-code generation has attracted significant interest as the visual elements of charts are more complex. This task challenges MLLMs to generate code that accurately reproduces a given chart or visual representation. Recent works Zhang et al. (2024d); Zhao et al. (2025) like ChartMimic Shi et al. (2024) evaluate the reasoning ability of MLLMs in this context. Similarly, Plot2Code Wu et al. (2024) and ChartX Xia et al. (2024) also evaluate MLLMs code generation ability, especially for text and data reproducibility. To the best of our knowledge, no dedicated research has focused on solving the chart-to-code generation problem. Our work is the first to attempt to address this challenge.
@@ 3.1 Direct Chart-to-code Generation
Despite the availability of many chart-reasoning instruction-tuning datasets, there is a notable lack of datasets specifically for chart-to-code tasks. Compared to chart reasoning data, chart-to-code data have the following distinct characteristics: (1) One-to-One Mapping: Unlike chart reasoning datasets, which could derive multiple question-answer pairs from a single chart, chart-to-code datasets require a one-to-one correspondence, demanding a large number of chart images for training. (2) Diversity Reflect on Charts: Unlike the diversity of chart reasoning data, which can be reflected in instructions, the diversity of chart-to-code data primarily lies in the variety of chart types and structures. (3) Syntax Constraints: Unlike the flexible natural language answers in chart reasoning tasks, the output code must strictly adhere to programming syntax to ensure executability.
@@
Specifically, we generate chart-to-code data through the following steps: First, we prompt the LLM to generate keywords within a specific domain and guide it to generate simulated data related to these keywords. Then, to ensure the diversity of chart types, we identify 27 commonly used chart types and manually write 79 template codes for each as in-context demonstrations. These template codes contain almost all common chart formats. We further provide available functions such as plt.text() and parameters such as hatch=’/’ to encourage the generation of more diverse functions and parameters, resulting in the chart structures more diversely. To enhance the generality of generated code, LLMs are encouraged to use standard libraries such as Matplotlib and Seaborn. Additionally, we explicitly define all parameters within the code itself, eliminating the need for external files such as CSVs. This ensures that the code can be executed directly and accurately to represent the chart details. The final step involved executing the generated code to produce the corresponding chart. We utilize the above process to yield 200k code snippets for charts. After executing the code and filtering out problematic charts, such as those with excessive pixels or ticks, we construct a high-quality dataset of 160k diverse chart-to-code pairs. These pairs are formatted as multimodal instruction-tuning samples in the unified structure of <chart, instruction, code>.
@@ 3.3 Dataset Analysis
Chart2Code-160k dataset provides three key advantages: (1) The First Large-Scale Dataset: It contains 160k data pairs for instruction tuning, significantly surpassing the size of previous datasets. (2) Diverse Chart Structures and Types: It includes 27 different chart types, with diverse structures enabled by a wide variety of functions and parameters in the code. (3) Syntactically Correct and Executable Code: All corresponding code is syntactically correct and executable, with explicitly defined parameters that ensure precise alignment between chart structures and code representations.
The comparisons of Chart2Code-160k with relevant chart-to-code datasets are listed in Table 1. To ensure data quality, we randomly sample around 1k instances and evaluate the quality of the chart images manually during the dataset construction period. Given the strong generation capabilities of LLM, we reckon the generated charts are suitable for training purposes. Furthermore, to quantitatively evaluate the chart quality, we also sample 8k data pairs (5% of the total) from Chart2Code-160k and utilize gpt-4o-2024-08-06 to evaluate them on four criteria: Aesthetics, Readability, Reproducibility, and Data Presentation Simplicity. The results in Table 2 show that the overall scores are broadly the same as real-world charts in ChartMimic. The detailed prompt is in the Figure 8. Our proposed Chart2Code-160k fills the gap between chart and code, equipping the model with advanced capabilities for downstream chart tasks.
Dataset Source Chart Quality
Mean $\mu$ SD $\sigma$
Chart2Code-160k sample Generated 77.32 4.04
ChartMimic testmini Real-world 78.96 3.96
Table 2: Quantitative evaluation of the chart quality, comparing with real-world charts. SD is the abbreviation version for standard deviation.
@@ 4 ChartCoder Model
After constructing the Chart2Code-160k, we aimed to leverage the data to enhance the capacities of MLLMs to generate code from charts. Unlike previous methods that rely on general LLMs with a low proportion of code in their training corpus, we pioneer the use of Code LLMs to enhance the coding abilities of MLLMs from scratch.
@@ 4.2 Model Training
Chart-to-text Alignment. The alignment process aims to endow the model with chart structure perception capability. In this stage, we freeze the language and vision encoder models and pre-train the vision-language connector Liu et al. (2023c). We collect and filter public chart corpora for alignment, which contains multiple tasks like chart caption and chart-to-table. Specifically, we use the following corpora: (1) UniChart Masry et al. (2023), (2) Chart-to-Text Kantharaj et al. (2022), (3) SciCap Hsu et al. (2021), and (4) SciCap+ Yang et al. (2024). Additionally, we incorporate the LLaVA pre-training dataset Liu et al. (2023c) and our proposed Chart2Code-160k to achieve a more balanced coverage of concepts.
@@ 5.2 Main Results
Table 5 shows the low-level results. ChartCoder even slightly outperforms GPT-4o in layout, type and color, highlighting the diversity of our proposed Chart2Code-160k dataset. However, the text score of the ChartCoder is lower than GPT-4o, which is similar to the results of high-level scores. We believe this is due to the lack of specialized chart OCR-oriented training for our model. Nevertheless, our text accuracy still surpasses that of open-source models, indicating the effectiveness of our proposed ChartCoder model and Chart2Code-160k dataset. We further present some case studies on ChartMimic and compare ChartCoder with existing MLLMs. The results are shown in Figure 3, the outputs of ChartCoder are much more similar to the ground truth chart than open-source models.
@@ 5.3 Ablation Study
We perform extensive ablation experiments to validate the effectiveness of our proposed model and dataset. We divide the ablation study into three parts, and the results are shown in Table 6. (1) Code or general LLMs. To investigate whether employing Code LLMs as language backbone provides specific advantages in chart-to-code tasks and identify the nature of these potential benefits, we replace the Code LLM, DeepSeek Coder 6.7B, with general LLM, DeepSeek LLM 7B Bi et al. (2024a), maintaining the same two-stage training procedures. The result shows that compared with general LLM, utilizing code LLM as the language backbone could significantly improve the execution rate, as well as the low-level and high-level scores. We further analyze the types of errors in the code that failed to execute and find that utilizing code LLMs significantly reduces syntax errors like missing closing quotation marks and type errors like incorrect argument type. (2) Resolution of vision encoders. Previous studies have indicated that performance on chart understanding tasks is resolution-dependent, with lower resolutions negatively impacting model performance Liu et al. (2024). To verify whether resolution affects chart-to-code tasks, we replace SigLIP-384 with CLIP-336 and maintain the other setting. The result shows that the resolution of the vision encoder generally does not affect the output code execution rate but slightly influences the high-level chart similarity. Through our analysis, we find that, similar to the challenges in chart understanding, this issue is caused by the negative impact of low resolution on the recognition of text and special symbols. However, as we utilize the Any Resolution strategy, this impact has been reduced significantly.
(3) Dataset effectiveness. We design two scenarios to illustrate our proposed Chart2Code-160k dataset. Firstly, to evaluate our proposed SoT method to emphasize the critical information in the chart, we remove the 50k step-by-step generation data and train the model using only the direct generation data. The result shows it influences the low-level and high-level scores notably, especially in text content and data, which shows the role of emphasising critical information. Secondly, we select Qwen2-VL-7B as the baseline of open-source MLLM and directly fine-tune it on our proposed Chart2Code-160k datasets. The result illustrates that after fine-tuning, the performance improves significantly on all the metrics, demonstrating the effectiveness of Chart2Code-160k.
Model Image Image+Code $\triangle$
MiniCPM-Llama3-V2.5 0.76 0.81 6.5%
InternVL2-8B 0.79 0.82 3.8%
Table 7: Comparison of the impact of using code as auxiliary contexts on the MMC True/False task.